Find a basis for the given subspace W of ?4 that includes the vectors (a,b,c,d) where c = a-b and d = a+b
Answer #1

that Given, а змb stra of te (a,b,c,d) includes the vectors where Ce a-b, d = atb . iie w= { Caibic,d) Eme 1 c-a-b, d=a+b} ht